### v1

@@ 3.2 Dataset Selection and Preprocessing
We base our study on the Penn Machine Learning Benchmark (PMLB) repository (olson2017pmlb), a curated collection of 450 datasets that has become a widely used benchmark suite, and is frequently used in studies on interpretable models, especially on SR (la2021contemporary; de2024srbench++). PMLB contains 179 classification datasets and 271 regression datasets. Not all of these are suitable for our purposes, so we apply a sequence of filtering and preprocessing steps summarized in Figure 1, where C and R denote the number of classification and regression datasets retained after each step. Orange boxes indicate filtering (dataset removal), while yellow boxes denote transformations applied to the datasets.
@@
Since most interpretable methods do not natively handle missing values, and imputation can distort feature distributions in a method-dependent way, we rely on simple deletion. First, we drop columns with more than 30% missing values, retaining only those datasets where at least half of the original columns remain. Next, we remove rows containing any missing values, keeping only datasets that preserve at least 50% of the original rows. Datasets requiring more extensive deletion are discarded, as they would be too heavily altered and could bias the results.
@@ 3.3 Hyperparameter Optimization
When available, the search space was initialized based on the hyperparameter ranges suggested in the papers introducing the corresponding methods. Hyperparameter ranges were further constrained to promote smaller and more comprehensible models, in line with interpretability objectives. For example, for Decision Trees, the maximum tree depth was limited to 6, as deeper trees would typically result in hundreds or thousands of nodes, making them difficult to interpret meaningfully. The full list of hyperparameters and their corresponding search ranges used in our experiments is reported in Table 3 in Appendix B. The hyperparameter optimization objective was the maximization of the performance measures described in Section 4.3.
Hyperparameter search was conducted using the Tree-structured Parzen Estimator (TPE) algorithm (bergstra2011algorithms), as implemented in the Optuna Python library (akiba2019optuna). This Bayesian optimization method allows for a more efficient exploration of the hyperparameter space compared to random search, which is particularly advantageous for high-dimensional configurations in some of the tested models. We executed at most 50 optimization trials per setting, limiting the total optimization time to 6000 seconds, corresponding to an average of approximately 2 minutes per trial. For models with higher computational demands, such as SR and GOSDT, this time constraint was explicitly enforced to ensure a feasible evaluation.
@@ 4.2 In-sample and out-of-sample generalization
    $\frac{l}{f}\alpha_{\mathcal{D}}\leq\alpha\leq\frac{u}{f}\alpha_{\mathcal{D}},$
@@ 4.3 Performance Metrics
$\text{F1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
We also computed several other standard metrics, such as Mean Squared Error (MSE) and Mean Absolute Error (MAE) for regression, as well as precision, recall, and accuracy for classification. Because these yielded results highly correlated with our primary metrics, we focus our discussion on $R^{2}$ and F1 to maintain clarity and avoid redundancy. Detailed results for all other metrics are provided in the supplementary material.
5 Model Evaluation
5.1 Dataset Metrics for Stratification
To better understand how IML models behave across heterogeneous problems, we characterize each dataset $\mathcal{D}$ using a small set of complexity descriptors that were proposed for classification (lorena_how_2020) and regression (maciel_measuring_2016). These descriptors summarize size, sparsity, class/target structure, and local neighborhood overlap, and are subsequently used to stratify performance across the PMLB benchmark.
@@ 5.1 Dataset Metrics for Stratification
  • Average number of PCA dimensions per point (T3):
    $\text{T3}=d^{\prime}/N\,$,
    where $d^{\prime}$ is the number of principal components explaining $95\%$ of the variance in $\mathbf{X}$.
These indicators jointly describe the scale and effective sparsity of a dataset. Larger values of $N$ typically stabilize empirical performance estimates, whereas small-sample regimes increase sensitivity to noise and model specification. High dimensionality ($d\gg 1$) can favor models with built-in regularization or feature selection while penalizing methods relying on local neighborhoods. The ratios T2 and T3 explicitly capture sparsity by relating dimensionality to sample size: large values indicate that data occupy a high-dimensional space relative to the number of observations, a regime in which decision boundaries are poorly supported and both parameter estimation and generalization become more challenging. In particular, T3 approximates intrinsic dimensionality after removing linear correlations, highlighting cases where many effective degrees of freedom must be learned from limited data.
@@
We characterize class separability, local overlap, and class balance using the following indicators:
@@
    $r_{j}=\frac{\sum_{c\in\mathcal{C}}p_{c}(\mu_{cj}-\mu_{j})^{2}}{\sum_{c\in\mathcal{C}}p_{c}\sigma_{cj}^{2}}\,$
@@
    $\text{N3}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\bigl(y_{i}^{\text{NN}}\neq y_{i}\bigr)\in[0,1].$
@@
    $\text{E5}=-\frac{1}{N}\sum_{i=1}^{N}\left[p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i})\right]\in[0,1]\,$
@@
    $p_{i}=\frac{1}{5}\sum_{j\in\mathcal{N}_{5}(i)}\mathbb{I}(y_{j}\neq y_{i})\,$,
@@
    $\text{C1}_{\text{c}}=-\frac{1}{\log C}\sum_{c\in\mathcal{C}}p_{c}\log p_{c}\in[0,1].$
@@
We characterize functional simplicity, linearity, and local smoothness using the following indicators:
@@
    $\text{C1}_{\text{r}}=\max_{j=1,\dots,d}|\rho(x_{j},y)|\in[0,1].$
@@
    $\text{L1}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-f(\mathbf{x}_{i})|,$
@@
    $\text{S3}=\frac{1}{N}\sum_{i=1}^{N}\bigl(\widehat{y}_{i}^{\text{NN}}-y_{i}\bigr)^{2},$
@@
Figure 3: Correlation plot of the metrics for classification datasets computed using the Spearman correlation.
Figure 4: Correlation plot of the metrics for regression datasets computed using the Spearman correlation.
The set of metrics considered, summarized in Table 2, spans multiple dimensions of dataset complexity. Their complementarity is assessed by examining pairwise Spearman correlations, reported in Figures 3 and 4. The predominantly low correlations across most pairs suggest these descriptors provide complementary information, justifying their joint use in our stratification. Notable exceptions include the expected structural dependencies between T2 and T3, as well as their natural correlation with $N$ and $d$. In classification, we observe a moderate correlation between N3 and E5, as both metrics probe local neighborhood purity. Similarly, for regression, $\text{C1}_{\text{r}}$ and L1 show a distinct correlation, reflecting their shared sensitivity to the linearity of the feature-target relationship.
In the following, we use these metrics to partition the PMLB benchmark into strata and analyze whether the relative performance of interpretable methods remains stable across these regions or changes substantially under more complex data conditions.
@@
Table 2: Complexity Metrics for Classification and Regression.
5.2 Model Ranking
For each dataset $\mathcal{D}$, models were ranked according to their predictive performance, with a rank of 1 assigned to the method exhibiting the best performance. We then calculated the average rank for each model either over the entire dataset collection or stratified by specific evaluation metrics introduced in Section 5.1. To formally compare the methods and assess the statistical significance of the observed differences in their average ranks, we employed the Friedman test (friedman1937use). This non-parametric test is appropriate for comparing multiple models across multiple datasets because it operates on ranks rather than raw performance scores. This is crucial as performance metrics across diverse datasets often violate the normality assumption and homoscedasticity required by parametric tests like ANOVA. If the null hypothesis (that all models perform equally well) was rejected by the Friedman test, we proceeded with the Nemenyi post-hoc procedure (nemenyi1963distribution) using a significance level of $\alpha=0.05$. The Nemenyi test determines the critical difference required for two average ranks to be considered statistically distinct, allowing us to identify groups of methods that are statistically indistinguishable from one another in terms of performance.
6 Regression Results
In this section, we present the results of our experiments for the regression task. Our analysis begins with an evaluation of the methods’ relative performance aggregated across the entire dataset collection in the in-sample setting. The results are then analyzed in greater detail by stratifying them according to the dataset metrics introduced in Section 5.1. Finally, we compare the methods’ performance in the Out-of-Sample configuration and examine their computational training time.
6.1 Regression Performance (In-Sample)
@@
Figure 5: Ranking comparison for the regression tasks between in-sample and out-of-sample settings.
Figure 5 shows the relative ranking of the regression methods based on the $R^{2}$ metric. We introduce a modified version of the critical-difference diagram in which brackets denote groups of models whose rankings are not statistically distinguishable, indicating comparable predictive performance. Models are ordered from bottom to top according to their average rank, from better to worse.
Figure 6: Comparison of model complexity, measured by the number of non-zero linear terms, between Linear Regression, LASSO, and Polynomial Regression with LASSO.
For the in-sample setting (left-hand side), we observe three clearly defined groups of methods. The EBM model outperforms the others by a substantial margin. The second group includes Polynomial Regression with LASSO regularization, IGANN, and Symbolic Regression. Notably, Symbolic Regression, typically valued for producing highly interpretable mathematical expressions, achieves performance comparable to that of GAMs, which are typically more complex to interpret. The final group comprises the remaining models, primarily linear regressors, Decision Trees, and k-NN, which generally exhibit lower performance.
Quantifying interpretability is generally architecture-specific, precluding a direct comparison across heterogeneous models (e.g., k-NN versus Symbolic Regression). However, within the family of linear models, we can structurally compare interpretability by analyzing the sparsity of the resulting equations. To quantify this sparsity, we apply a numerical threshold, treating any coefficient with an absolute value below $10^{-6}$ as zero. Figure 6 illustrates the distribution of the remaining non-zero coefficients, showing that LASSO effectively minimizes model complexity compared to standard Linear Regression. In contrast, PR+LASSO exhibits a marked increase in the number of terms due to the expansion of the feature space. Beyond the sheer number of terms, the inclusion of higher-order transformations and interaction effects makes Polynomial Regression inherently less transparent than purely linear models. At the same time, this added structural complexity enables the model to represent non-linear relationships that are fundamentally inaccessible to linear formulations. Therefore, PR+LASSO exemplifies a clear performance–interpretability trade-off: increased expressive power can yield improved predictive accuracy, but results in symbolic representations that are more elaborate and require greater effort to interpret.
6.1.1 Stratified Results for Regression

### v2

@@ 3.2 Dataset Selection and Preprocessing
We base our study on the Penn Machine Learning Benchmark (PMLB) repository [olson2017pmlb], a curated collection of 450 datasets that has become a widely used benchmark suite, and is frequently used in studies on interpretable models, especially on SR [la2021contemporary, de2024srbench++]. PMLB contains 179 classification datasets and 271 regression datasets. Not all of these are suitable for our purposes, so we apply a sequence of filtering and preprocessing steps summarized in Figure 1, where C and R denote the number of classification and regression datasets retained after each step. Orange boxes indicate filtering (dataset removal), while yellow boxes denote transformations applied to the datasets.
@@
Since most interpretable methods do not natively handle missing values, and imputation can distort feature distributions in a method-dependent way, we rely on simple deletion. First, we drop columns with more than 30% missing values, retaining only those datasets where at least half of the original columns remain. Next, we remove rows containing any missing values, keeping only datasets that preserve at least 50% of the original rows. Datasets requiring more extensive deletion are discarded, as they would be too heavily altered and could bias the results. Notably, under these criteria, only a single classification dataset was excluded. To further ensure that this deletion strategy does not introduce bias, Appendix C provides a detailed quantification of the missing data, demonstrating that the proportion of removed cells in the retained datasets is negligible, with most datasets containing no missing values at all.
@@ 3.3 Hyperparameter Optimization
When available, the search space was initialized based on the hyperparameter ranges suggested in the papers introducing the corresponding methods. Hyperparameter ranges were further constrained to promote smaller and more comprehensible models, in line with interpretability objectives. For example, for Decision Trees, the maximum tree depth was limited to 6, as deeper trees would typically result in hundreds or thousands of nodes, making them difficult to interpret meaningfully. The full list of hyperparameters and their corresponding search ranges used in our experiments is reported in Table D1 in Appendix D. The hyperparameter optimization objective was the maximization of the performance measures described in Section 4.3.
Hyperparameter search was conducted using the Tree-structured Parzen Estimator (TPE) algorithm [bergstra2011algorithms], as implemented in the Optuna Python library [akiba2019optuna]. This Bayesian optimization method allows for a more efficient exploration of the hyperparameter space compared to random search, which is particularly advantageous for high-dimensional configurations in some of the tested models. We executed at most 50 optimization trials per setting, limiting the total optimization time to 6000 seconds, corresponding to an average of approximately 2 minutes per trial. For models with higher computational demands, such as SR, this time constraint was explicitly enforced to ensure a feasible evaluation. While unconstrained search times might yield marginal improvements for computationally heavy methods, our time budget was strictly enforced to reflect realistic operational constraints and to make such a large-scale evaluation possible.
@@ 4.2 In-sample and out-of-sample generalization
    $\frac{l}{f}\alpha_{\mathcal{D}}\leq\alpha\leq\frac{u}{f}\alpha_{\mathcal{D}},$ (1)
@@ 4.3 Performance Metrics
$\text{F1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$ (4)
We also computed several other standard metrics, such as Mean Squared Error (MSE) and Mean Absolute Error (MAE) for regression, as well as precision, recall, accuracy, and Brier score for classification. Because these yielded results highly correlated with our primary metrics, we focus our discussion on $R^{2}$ and F1 to maintain clarity and avoid redundancy. Detailed results for all other metrics are provided in the supplementary material.
4.4 Meta-Features for Dataset Stratification and Algorithm Selection
To better understand how IML models behave across heterogeneous problems, we frame our dataset characterization within the broader context of meta-learning. We characterize each dataset $\mathcal{D}$ using a small set of complexity descriptors, or meta-features, that were proposed for classification [lorena_how_2020] and regression [maciel_measuring_2016]. These descriptors summarize size, sparsity, class/target structure, and local neighborhood overlap, and are subsequently used to stratify performance across the PMLB benchmark.
@@ 4.4 Meta-Features for Dataset Stratification and Algorithm Selection
  • Average number of PCA dimensions per point (T3): $\text{T3}=d^{\prime}/N\,$,\; where $d^{\prime}$ is the number of principal components explaining $95\%$ of the variance in $\mathbf{X}$.
These meta-features jointly describe the scale and effective sparsity of a dataset. Larger values of $N$ typically stabilize empirical performance estimates, whereas small-sample regimes increase sensitivity to noise and model specification. High dimensionality ($d\gg 1$) can favor models with built-in regularization or feature selection while penalizing methods relying on local neighborhoods. The ratios T2 and T3 explicitly capture sparsity by relating dimensionality to sample size: large values indicate that data occupy a high-dimensional space relative to the number of observations, a regime in which decision boundaries are poorly supported and both parameter estimation and generalization become more challenging. In particular, T3 approximates intrinsic dimensionality after removing linear correlations, highlighting cases where many effective degrees of freedom must be learned from limited data.
@@
We characterize class separability, local overlap, and class balance using the following complexity measures:
@@
    $r_{j}=\frac{\sum_{c\in\mathcal{C}}p_{c}(\mu_{cj}-\mu_{j})^{2}}{\sum_{c\in\mathcal{C}}p_{c}\sigma_{cj}^{2}}\,$ (6)
@@
    $\text{N3}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\bigl(y_{i}^{\text{NN}}\neq y_{i}\bigr)\in[0,1].$ (7)
@@
    $\text{E5}=-\frac{1}{N}\sum_{i=1}^{N}\left[p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i})\right]\in[0,1]\,$ (8)
@@
    $p_{i}=\frac{1}{5}\sum_{j\in\mathcal{N}_{5}(i)}\mathbb{I}(y_{j}\neq y_{i})\,$, (9)
@@
    $\text{C1}_{\text{c}}=-\frac{1}{\log C}\sum_{c\in\mathcal{C}}p_{c}\log p_{c}\in[0,1].$ (10)
@@
We characterize functional simplicity, linearity, and local smoothness using the following complexity measures:
@@
    $\text{C1}_{\text{r}}=\max_{j=1,\dots,d}|\rho(x_{j},y)|\in[0,1].$ (11)
@@
    $\text{L1}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-f(\mathbf{x}_{i})|,$ (12)
@@
    $\text{S3}=\frac{1}{N}\sum_{i=1}^{N}\bigl(\widehat{y}_{i}^{\text{NN}}-y_{i}\bigr)^{2},$ (13)
@@
((a)) Regression
((b)) Classification
Figure 3: Correlation plot of the meta-features for (a) regression and (b) classification datasets computed using the Spearman correlation.
@@
Table 2: Meta-Features for Classification and Regression.
The set of meta-features considered, summarized in Table 2, spans multiple dimensions of dataset complexity. Their complementarity is assessed by examining pairwise Spearman correlations, reported in Figures 3(a) and 3(b). The predominantly low correlations across most pairs suggest these descriptors provide complementary information, justifying their joint use in our stratification. Notable exceptions include the expected structural dependencies between T2 and T3, as well as their natural correlation with $N$ and $d$. In classification, we observe a moderate correlation between N3 and E5, as both metrics probe local neighborhood purity. Similarly, for regression, $\text{C1}_{\text{r}}$ and L1 show a distinct correlation, reflecting their shared sensitivity to the linearity of the feature-target relationship.
In the following, we use these meta-features to partition the PMLB benchmark into strata and analyze whether the relative performance of interpretable methods remains stable across these regions or changes substantially under more complex data conditions. Building on this stratification, we then operationalize these metrics by training meta-models to predict both pairwise algorithm superiority and inherent dataset difficulty.
4.5 Evaluation of Interpretability
As discussed, measuring true cognitive notion of model comprehensibility often relies on subjective human evaluations, which lack universal consensus and are nearly impossible to standardize across diverse model architectures. Since we cannot directly compare interpretability across heterogeneous model classes, this study primarily quantifies predictive performance, generalization, and computational efficiency.
To enable a more controlled comparison, we instead assess interpretability within homogeneous model families. In this context, we operationalize interpretability through structural parsimony. As detailed in Section 5.5, we introduce objective proxies for model complexity: for linear models, interpretability is quantified by the number of non-zero coefficients, while for tree-based methods, we consider the total number of nodes. These measures provide consistent, quantifiable bounds on the cognitive effort required to understand the models.
5 Results
((a)) Regression datasets
((b)) Classification datasets
Figure 4: (a) Regression and (b) classification performance profile across datasets (sorted along the x-axis by the median performance).
((a)) Regression
((b)) Classification
Figure 5: Predictive accuracy distribution of the methods on all the datasets for (a) regression ($R^{2}$ scores) and (b) classification (F1 scores).
Before comparing the specific methods, we assess the performance variability across the dataset collection to ensure that the benchmark provides a sufficiently diverse and challenging evaluation testbed.
By analyzing the median predictive performance ($R^{2}$ for regression, Figure 4(a), and F1 score for classification, Figure 4(b)) on each dataset alongside the range between the best and worst-performing models, we observe a wide spectrum of dataset difficulties. For both tasks, the benchmark includes datasets where models achieve near-perfect scores (approaching 1) and others where models struggle significantly. Notably, the performance range in regression is quite wide. In classification, while the range is less pronounced, there remains a non-negligible gap between the best and worst models. This indicates that the benchmark is not saturated; the models have not converged on an intrinsic performance ceiling, ensuring that the datasets are suitably complex for the comparative analysis.
Moving to the performance of the individual methods (Figure 5(a) for regression, and Figure 5(b) for classification), the score distributions reveal distinct behaviors between the two tasks. In regression, we observe a considerable difference among the methods. EBM and Symbolic Regression achieve the highest overall performance, followed by IGANN and PR+LASSO. k-NN exhibits high variability, followed by Decision Trees, while the linear models (Linear Regression, LASSO, and GLM) perform similarly at the lower end of the spectrum.
In classification, the distinction between models is much less clear. EBM maintains a slight performance advantage, but the remaining models exhibit highly overlapping performance distributions. The only notable exception is Naive Bayes, which performs poorly across the majority of the datasets.
5.1 Method Rankings and Pairwise Comparisons
((a)) In-sample vs. Out-Of-Sample rankings (regression)
((b)) In-sample vs. Out-Of-Sample rankings (classification)
@@ 5.1 Method Rankings and Pairwise Comparisons
Figure 6: Ranking comparison for (a) the regression and (b) the classification tasks between in-sample and out-of-sample settings.
((a)) Regression
((b)) Classification
Figure 7: Pairwise comparison between methods in the in-sample setting for (a) regression and (b) classification.
To rigorously establish a performance hierarchy, we evaluated the statistical significance of the model rankings. Specifically, we utilized the Friedman test, a non-parametric alternative to repeated-measures ANOVA used to detect overall differences across multiple datasets, paired with the Nemenyi post-hoc test, which identifies specific pairwise differences by comparing mean ranks. This approach follows established best practices for comparing machine learning algorithms [demvsar2006statistical]. In our critical difference visualizations, brackets connect methods whose performances are not statistically distinguishable. Furthermore, to compare the in-sample and out-of-sample settings, we applied the Wilcoxon Signed-Rank Test, a non-parametric method used to assess whether the median difference between paired observations is statistically significant [demvsar2006statistical]. To control the family-wise error rate across repeated measures, we incorporated a Holm correction. Significant trends in these ranks are denoted by solid lines.
For regression in the in-sample setting (Figure 6(a)), three distinct groups emerge. EBM isolates itself as the top-performing method. The second tier comprises SR, IGANN, and PR+LASSO, while the final tier includes the remaining algorithms. In the out-of-sample setting, the overall hierarchy remains largely intact, though PR+LASSO suffers a relative decrease in rank, together with the decision trees. EBM, SR, and IGANN consistently remain the top performers, and regularized linear models like LASSO tend to be more robust in this setting. Conversely, the classification rankings, in Figure 6(b), are highly overlapping, making it difficult to distinguish clear, statistically significant tiers among the majority of the algorithms.
This trend is further corroborated by our analysis of the pairwise win probabilities. We calculated the probability of one method outperforming another, employing the Wilcoxon signed-rank test with a Holm correction to identify statistically significant pairs. Crucially, unlike previous rank-based evaluations, this test assesses raw performance values; thus, marginal differences may lack statistical significance even if one method consistently outranks another. In the regression task (Figure 7(a)), the performance hierarchy is distinct, with differences being almost universally significant. However, there are two notable exceptions where no statistically significant difference is observed: between EBM and SR, as well as between LR and LASSO. These exceptions are particularly relevant as they demonstrate that (i) SR, despite being constrained to provide analytical and highly interpretable models, yields highly competitive results, and (ii) the application of LASSO regularization does not compromise predictive performance relative to standard linear regression.
In classification (Figure 7(b)), the situation is drastically different. Aside from EBM’s dominance and Naive Bayes’ underperformance, a definitive hierarchy among the remaining methods cannot be established, corroborating the ranking analysis.
These findings highlight an intrinsic difference between the regression and classification tasks. While the regression analysis reveals distinct behavioral differences among the methods, the classification results exhibit a less pronounced performance hierarchy. In the subsequent sections, we deepen our analysis by characterizing the datasets using the meta-features introduced in Section 4.4. Guided by our initial findings, the regression analysis will focus on the relationship between these meta-features and model performance. Conversely, the classification analysis will investigate whether meta-features can establish a clearer hierarchy, or if underlying patterns and method clusters exist that were missed in the aggregated analysis.
5.2 Meta-Features Effects on Regressors
((a)) Regression
((b)) Classification
Figure 8: Spearman correlation between dataset meta-features and (a) regressor or (b) classifier performance.